# Tuning Language Models for Robust Prediction of Diverse User Behaviors

### v1

@@ Abstract
Predicting user behavior is essential for intelligent assistant services, yet deep learning models often struggle to capture long-tailed behaviors. Large language models (LLMs), with their pretraining on vast corpora containing rich behavioral knowledge, offer promise. However, existing fine-tuning approaches tend to overfit to frequent “anchor” behaviors, reducing their ability to predict less common “tail” behaviors. In this paper, we introduce BehaviorLM, a progressive fine-tuning approach that addresses this issue. In the first stage, LLMs are fine-tuned on anchor behaviors while preserving general behavioral knowledge. In the second stage, fine-tuning uses a balanced subset of all behaviors based on sample difficulty to improve tail behavior predictions without sacrificing anchor performance. Experimental results on two real-world datasets demonstrate that BehaviorLM robustly predicts both anchor and tail behaviors and effectively leverages LLM behavioral knowledge to master tail behavior prediction with few-shot examples.
1 Introduction
@@
Most people lead routine lives shaped by habits, but also display short-term bursts of context-driven activities. Accurately predicting users’ next behavior is essential for intelligent assistant services across web platforms and smart devices Chung and Lee (2018); Tulshan and Dhage (2019); Savcisens et al. (2023); Rychalska et al. (2023).
@@
With the accumulation of user behavior data, deep learning has replaced rule-based methods Zhang and Dai (2018); Zhang et al. (2019); Li et al. (2022), with transformer-based models excelling at capturing sequential patterns Kang and McAuley (2018); Pu et al. (2018); Sun et al. (2019); Savcisens et al. (2023). However, their reliance on extensive training data leads to high costs. In this context, Large language models (LLMs) Brown (2020); Zhao et al. (2023) offer a promising alternative by leveraging pretrained knowledge of human behavior, showing potential in interpreting user intent and generating accurate predictions Wu et al. (2024); Zhao et al. (2024).
@@
Recent works adapt LLMs for behavior prediction by converting behavioral sequences into text and fine-tuning them to predict user behaviors Li et al. (2023); Bao et al. (2023); Liao et al. (2024); Kim et al. (2024). Early methods Bao et al. (2023) combined general instruction tuning with lightweight behavior fine-tuning, and subsequent research Liao et al. (2024) enhanced LLMs with embeddings from traditional models. While expressing user behaviors as text appears intuitive Li et al. (2023), the long-tailed distribution of behaviors presents a significant challenge Liu et al. (2024): a user’s daily life is predominantly characterized by a small subset of frequent behaviors serving as “anchors”, while other behaviors occur far less commonly. Recent work Liu et al. (2024) explores hybrid methods, combining LLMs’ profiling capabilities with traditional models to address this issue. However, capturing long-tailed distributions remains a persistent challenge for LLMs. Our empirical observations reveal that while an 8B-parameter LLM fine-tuned on behavior data outperforms untuned GPT-4 on anchor behaviors, it still underperforms on tail behaviors. This highlights a fundamental limitation: fine-tuning alone does not uniformly enhance prediction capabilities across diverse behaviors.
@@
In this paper, we empirically investigate LLMs’ uneven performance across behavior types and make a surprising discovery: an LLM fine-tuned only on anchor behaviors can still generalize well to tail behaviors in a zero-shot manner, even better than full data fine-tuning. Motivated by this finding, we propose BehaviorLM, a two-stage fine-tuning progressive framework designed to better model long-tailed user behaviors. In the first stage, we fine-tune the LLM using anchor behaviors and a general conversational corpus in a multi-task setup to preserve general behavioral knowledge and prevent bias. In the second stage, we fine-tune the model on a balanced subset of all behaviors, selected by sample difficulty, to improve tail behavior predictions without compromising anchor performance. Unlike existing methods that rely on auxiliary data such as embeddings from traditional deep learning-based models, increasing training complexity and limiting generalization, our approach leverages the intrinsic semantic knowledge of LLMs to overcome these challenges.
Our main contributions are as follows:
This paper introduce a novel progressive fine-tuning method for adapting LLM into the robust predictor of user daily behaviors, addressing the challenge of long-tailed user preferences.
@@
We propose a progressive fine-tuning method that effectively leverages LLMs’ general behavioral knowledge for accurate and efficient behavior prediction.
Extensive experiments on two real-world datasets show BehaviorLM achieves up to 27.4%/20.4% accuracy improvements on tail behaviors, offering 100 higher sample efficiency than traditional transformer models and enabling few-shot prediction with as few as 20 examples. Ablation studies further validate the design of our method.
2 Motivation
@@
2.1 Problem Formulation
Next behavior prediction refers to the task of predicting a user’s next behavior, , given a chronologically ordered sequence of their most recent historical behavioral events. Each event indicates that a specific behavioral event occurred at location and time . The behavioral event specifically refers to daily activities—such as exercise or gaming—rather than fine-grained actions like picking up a cup. The location indicates a semantic indicator such as "home" or "workplace." The time includes both date and hour information.
@@ 2.1 Problem Formulation
2.2 Investigating LLM Fine-tuning Performance on Diverse User Behavior Dataset
@@
In this subsection, we evaluate the behavior prediction performance of fine-tuning approaches for LLMs using a real-world dataset of 37 smartphone user behaviors (detailed in Section 4). As shown in Figure 1(a), there is a significant class-imbalance issue, with behaviors divided into anchor (occurrence >1%) and tail categories. Despite only 16 anchor behaviors, they represent over 97% of the data, with an average anchor-to-tail occurrence ratio of 42.44 (0.97/16 vs. 0.03/21) and a maximum exceeding 2,500 (43.6% vs. 0.0163%).
@@
To further analyze the semantic meaning and similarity of diverse behaviors from the LLM’s perspective, we use the pretrained Llama-8B v3.1 Dubey et al. (2024) to generate behavior embeddings and reduce them from 4096 to 2 dimensions using PCA Abdi and Williams (2010). As shown in Figure 1(b), anchor behaviors serve as semantic anchors in the latent space, with tail behaviors clustering around them. For example, watching videos, especially short ones, is highly frequent, while nearby behaviors like watching sports or animation occur less often and appeal to fewer users. Similarly, public transportation is a more common commuting behavior than hailing a taxi.
To investigate LLM behavior prediction performance, we fine-tune Llama-8B v3.1 on the previously described behavior dataset and compare average prediction accuracy for anchor and tail behaviors using an equal number of samples per behavior type (Figure 1(c)). For reference, GPT-4 (gpt-4o-2025-03-26) achieves accuracy scores of 0.45 for anchor behaviors and 0.33 for tail behaviors. In its base form, Llama-8B v3.1 performs worse than GPT-4 for both anchor and tail behaviors. While fine-tuning improves its performance on anchor behaviors, surpassing GPT-4, it continues to underperform on tail behaviors. GPT-4’s strong overall performance, achieved without behavior-specific fine-tuning, suggests current fine-tuning methods overly favor anchor behaviors at the expense of tail behaviors, likely due to the severe class imbalance shown in Figure 1(a).
@@
Motivated by these observations, we explored whether addressing class imbalance during fine-tuning could better utilize the LLM’s inherent understanding of user behaviors. We fine-tuned the model exclusively on anchor behavior data, omitting tail behaviors. Notably, this approach preserved strong performance on anchor behaviors while enabling robust zero-shot prediction for tail behaviors, achieving an accuracy of 0.39—significantly higher than the 0.29 accuracy from traditional fine-tuning. This suggests that focusing on anchor behaviors allows the LLM to capture fundamental user behavior patterns and leverage its general knowledge to generalize to semantically related, unseen behaviors. This insight underpins our progressive fine-tuning approach for BehaviorLM.
3 Method
3.1 The BehaviorLM Framework
3.1.1 Behavior Prediction as Language Modeling
@@
To frame the next behavior prediction task as a language modeling problem, we design a specific prompt and apply instruction fine-tuning to the LLM.
@@
Prompt design. We employ a text-only prompt to convert user behavior history and the next behavior (if provided) into text-based metadata suitable for instruction tuning. The prompt consists of five components:
1.
@@
Task definition: A concise description of the task, e.g., "Predict the user’s next behavior from the candidates.
2.
@@
Role-playing instruction: Defines the LLM’s role and task Shanahan et al. (2023), e.g., "You are a smart mobile assistant that can infer the user’s behavior preferences based on their history."
3.
@@
Historical behavior sequence (Input): Transforms each element in into a textual description, e.g., "(1, 16, home, Exercise), (1, 20, home, Gaming), (1, 20, home, Video)."
4.
@@
Candidate set (Input): Specifies all possible behaviors for prediction, e.g., "Weather Check, Music, Cycling, Reading."
5.
@@
This design transforms user behavior data into a format optimized for LLM instruction fine-tuning.
@@
Instruction Fine-Tuning LLMs. To quickly adapt an LLM for behavior prediction tasks, we use instruction fine-tuning Ouyang et al. (2022). The user behavior dataset is converted into a text-based instruction dataset , where and are natural language representations of inputs and outputs. The LLM is optimized with the next-token prediction loss:
where are the LLM parameters, is the -th token of , and represents tokens preceding . Moreover, we adopt a well-known parameter-efficient fine-tuning technique, i.e., LoRA Hu et al. (2021), to fine-tune the LLM more efficiently.
3.1.2 Progressive Fine-Tuning with Behavior Data
@@
As illustrated in Figure 1(a), user behavior data often exhibits a significant class imbalance, with many behaviors occurring infrequently, leading to a long-tailed distribution. When the task of predicting a user’s next behavior is reframed as generating the corresponding language token, this imbalance persists, causing poor performance on infrequent (tail) behaviors, as shown by the results of "Llama3.1-8B-FT" in Figure 1(b). Interestingly, we observe that fine-tuning an LLM on a subset of frequently occurring behaviors (anchor behaviors) enables it to act as a surprisingly effective zero-shot predictor for less frequent (tail) behaviors, often outperforming fine-tuning on the full dataset. Inspired by this, we propose a progressive fine-tuning strategy to enhance BehaviorLM’s ability to predict a diverse range of user behaviors.
@@
The proposed fine-tuning strategy consists of two progressive stages (Figure 2). In the first stage, referred to as A-Tuning, the LLM is fine-tuned using anchor behavior data to specialize in user behavior prediction while retaining its inherent rich knowledge of long-tailed behaviors. In the second stage, referred to as B-Tuning, the model is further fine-tuned on a class-balanced dataset covering all behaviors in a few-shot manner, enhancing its predictive capability for tail behaviors. The first stage helps the LLM become a specialist in predicting anchor behaviors, while the second stage transforms it into a generalist capable of predicting both anchor and tail behaviors.
3.2 Multitask-Enhanced Anchor Behavior Tuning (A-Tuning)

### v2

@@ Abstract.
Predicting user behavior is essential for intelligent assistant services, yet deep learning models often struggle to capture long-tailed behaviors. Large language models (LLMs), with their pretraining on vast corpora containing rich behavioral knowledge, offer promise. However, existing fine-tuning approaches tend to overfit to frequent “anchor” behaviors, reducing their ability to predict less common “tail” behaviors. In this paper, we introduce BehaviorLM, a progressive fine-tuning approach that addresses this issue. In the first stage, LLMs are fine-tuned on anchor behaviors while preserving general behavioral knowledge. In the second stage, fine-tuning uses a balanced subset of all behaviors based on sample difficulty to improve tail behavior predictions without sacrificing anchor performance. Experimental results on two real-world datasets demonstrate that BehaviorLM robustly predicts both anchor and tail behaviors and effectively leverages LLM behavioral knowledge to master tail behavior prediction with few-shot examples. Code and data are available at https://anonymous.4open.science/r/1015-new-E5F0
1. Introduction
@@
Most people lead routine lives, driven by behavioral habits, but also exhibit short-term bursts of activity influenced by specific contexts. The ability to predict users’ next behavior is crucial for modern intelligent assistant services, which span a wide range of applications, from web platforms to smart devices (Chung and Lee, 2018; Tulshan and Dhage, 2019; Savcisens et al., 2023; Rychalska et al., 2023).
@@
With the accumulation of empirical data on user behaviors, deep learning-based approaches have replaced traditional rule-based methods (Zhang and Dai, 2018; Zhang et al., 2019; Li et al., 2022) and become the mainstream solution in this field. Among them, transformer-based models (Kang and McAuley, 2018; Pu et al., 2018; Sun et al., 2019; Savcisens et al., 2023) excel at capturing transition patterns between sequential behaviors. However, sequential behavior modeling typically demands extensive training data, which incurs significant costs. In this context, large language models (LLMs) (Brown, 2020; Zhao et al., 2023) have emerged as an ideal choice for behavior prediction, as these pretrained models inherently encode knowledge about human behavior from their vast training corpora. They show promise in interpreting the underlying intent behind observed actions and generating corresponding predictions (Wu et al., 2024b; Zhao et al., 2024).
@@
To adapt LLMs for behavior prediction, current approaches typically convert behavioral sequence data into textual format and fine-tune LLMs to predict user behavior tokens (Li et al., 2023; Bao et al., 2023; Liao et al., 2024; Kim et al., 2024). Among them, an early attempt (Bao et al., 2023) combines general instruction tuning on conversation data with lightweight fine-tuning on behavior data. Building on this, subsequent research (Liao et al., 2024) enhances LLMs’ behavioral understanding by incorporating traditional models’ encoded embeddings into text instructions. While expressing user behaviors as text appears intuitive (Li et al., 2023), the long-tailed distribution of behaviors presents a significant challenge (Liu et al., 2024). Specifically, a user’s daily life is predominantly characterized by a small subset of frequent behaviors serving as “anchors”, while other behaviors occur far less commonly. Recent work (Liu et al., 2024) proposes combining LLMs’ capacity to characterize item profiles, particularly for long-tailed items, with traditional sequential recommendation modules. However, capturing this long-tailed distribution remains a persistent challenge for LLMs. Our empirical observations reveal that an 8B-parameter LLM, after behavior data fine-tuning, readily outperforms the more powerful but untuned GPT-4 on anchor behaviors, yet falls short on tail behaviors. This disparity suggests that fine-tuning alone does not uniformly enhance prediction capabilities across diverse behaviors.
@@
In this paper, we empirically investigate LLMs’ uneven performance across diverse behaviors after fine-tuning and, surprisingly, find that an LLM fine-tuned exclusively on anchor behaviors can effectively predict tail behaviors in a zero-shot manner. Inspired by this novel finding, we propose a progressive fine-tuning approach to better capture the long-tailed distribution of user behavior. In the first stage, to preserve general behavioral knowledge and prevent bias toward overrepresented behaviors, we fine-tune the LLM not only with anchor behaviors but also by integrating a general conversational corpus in a multi-task manner. This allows the LLM to specialize in predicting anchor behaviors while retaining the ability to generalize to unseen tail behaviors. In the second stage, we further fine-tune the LLM across all behavior types by selecting a balanced subset of data based on sample difficulty. This approach significantly enhances prediction accuracy for tail behaviors without compromising performance on anchor behaviors. This two-stage fine-tuning framework, which we call BehaviorLM, effectively leverages the LLM’s general behavioral knowledge for sample-efficient tuning, enabling robust predictions across both anchor and tail behaviors.
To summarize, our main contributions are as follows.
We introduce a novel approach for adapting LLMs to the human behavioral domain, addressing the longstanding challenge of capturing long-tailed user preferences.
@@
We propose a progressive LLM fine-tuning strategy, inspired by empirical findings, that effectively leverages the general behavioral knowledge stored in LLMs for behavior prediction.
Experimental results on two real-world datasets demonstrate the superiority of BehaviorLM over state-of-the-art baselines in behavior prediction. Notably, BehaviorLM achieves a 30.8%/22.5% improvement in prediction accuracy for tail behaviors, underscoring its ability to capture long-tailed preferences. In-depth analysis reveals that the behavioral knowledge stored in the LLM provides over a 100 improvement in sample efficiency compared to traditional transformer-based approaches, while also enabling few-shot ( 20) predictive capability for tail behaviors. Ablation studies further validate the rationale behind the design of our method.
2. Motivation
@@
2.1. Problem Formulation
Next behavior prediction refers to the task of predicting a user’s next behavior, , given a chronologically ordered sequence of their most recent historical behavioral events. Each event indicates that a specific behavioral event occurred at location and time . The behavioral event specifically refers to daily activities—such as exercise or gaming—rather than fine-grained actions like picking up a cup. The location indicates a semantic indicator such as ”home” or ”workplace.” The time includes both date and hour information.
@@ 2.1. Problem Formulation
To facilitate understanding, we summarize the key mathematical notations used in this paper in Table 1.
| Notations | Description |
| The set of all possible user behaviors (candidate set) | |
| The historical behavior sequence of a user | |
| The ground truth next behavior to be predicted | |
| A specific behavioral event tuple (location, time, behavior) | |
| The length of the historical behavior sequence | |
| The collected raw user behavior dataset | |
| The instruction tuning dataset converted from | |
| The subset of anchor behaviors (frequent behaviors) | |
| The subset of tail behaviors (infrequent behaviors) | |
| The auxiliary conversation dataset for multi-task learning | |
| Ratio controlling the size of auxiliary task data | |
| The learnable parameters of the LLM | |
| The reference model obtained from the A-Tuning stage | |
| The policy model being optimized in the B-Tuning stage | |
| Prediction confidence of the model for behavior | |
| The confusion-based penalty score | |
| The continuous difficulty coefficient for a sample | |
| Hyper-parameter balancing uncertainty and confusion penalty | |
| Hyper-parameter controlling deviation in DPO loss | |
| The Direct Preference Optimization (DPO) loss function |
@@
2.2. Investigating LLM Fine-tuning Performance on Diverse User Behavior Dataset
@@
In this subsection, we empirically investigate the behavior prediction performance of current fine-tuning approaches for LLMs. We utilize a real-world dataset recording 37 types of daily user behaviors on a smartphone (detailed in Section 4). As shown in Figure 1(a), there is a significant class-imbalance issue in terms of occurrence frequency among different behavior types. Without loss of generality, we divided these behaviors into two categories: anchor behaviors, with an occurrence frequency greater than 1%, and tail behaviors for the rest. Although there are only 16 types of anchor behaviors, they account for over 97% of the total data, with an average ratio between anchor and tail behaviors of approximately 42.44 (0.97/16 vs. 0.03/21), and the highest ratio exceeding 2,500 (43.6% vs. 0.0163%).
@@
To further analyze the semantic meaning and similarity of these diverse behaviors from the perspective of the LLM, we use a pretrained LLM, Llama-8B v3.1 (Dubey et al., 2024), to generate embedding vectors for the behaviors and visualize them in two-dimensional space (reduced from 4096 dimensions) using PCA (Abdi and Williams, 2010). As shown in Figure 1(b), anchor behaviors act as semantic anchors in the latent space, scattered across it, with several tail behaviors clustering around these anchor points. For example, watching videos, especially short videos, has become a highly frequent behavior in users’ daily lives, while other nearby behaviors in the figure, such as watching sports matches or animation, occur less frequently and are favored by fewer individuals. Similarly, public transportation is a more common commuting behavior than hailing a taxi for most people.
To investigate the behavior prediction performance of LLMs, we use Llama-8B v3.1 as the backbone model and fine-tune it on the behavior dataset described earlier. For evaluation, we select an equal number of samples for each behavior type and report the average prediction accuracy for anchor and tail behaviors, respectively, as shown in Figure 1(c). For comparison, we first evaluate the prediction accuracy of GPT-4 (version gpt-4o-2024-08-06), which yields accuracy values of 0.45 for anchor behaviors and 0.33 for tail behaviors. In its base form, LLaMA3.1-8B significantly underperforms GPT-4 in predicting both anchor and long-tail behaviors. After applying established fine-tuning practices (Bao et al., 2023; Liao et al., 2024), the model shows marked improvement on anchor behaviors, surpassing GPT-4’s performance. However, for long-tail behaviors, it continues to lag behind GPT-4’s capabilities. GPT-4’s superior comprehensive performance across all behavior types, achieved without behavior-specific fine-tuning, suggests that current fine-tuning approaches may disproportionately optimize for anchor behaviors at the expense of tail behaviors. We attribute this performance disparity to the severe class imbalance illustrated in Figure 1(a).
@@
Motivated by these observations, we investigated whether eliminating class imbalance during fine-tuning could better leverage the LLM’s inherent knowledge of user behaviors. We fine-tuned the LLM exclusively on anchor behavior data, excluding all tail behavior samples. As illustrated in the figure, this approach maintains strong performance on anchor behaviors. Remarkably, despite having no exposure to tail behaviors during fine-tuning, the model demonstrates robust zero-shot predictive capabilities for these behaviors, achieving an accuracy of 0.39—substantially higher than the 0.29 obtained through traditional fine-tuning. This finding suggests that training on anchor behaviors alone enables the LLM to develop a fundamental understanding of user behavior patterns and leverage its general knowledge to generalize to unseen, semantically related behaviors. This insight forms the foundation of our progressive fine-tuning approach for BehaviorLM.
3. Method
3.1. The BehaviorLM Framework
3.1.1. Behavior prediction as language modeling
@@
To transform the next behavior prediction task into a language modeling task, we first design the specific prompt and then adopt the instruction fine-tuning technique for LLMs.
@@
Prompt design. We adopt a text-only prompt design that represents the behavioral history and the next behavior (if given) through textual metadata embedded within the prompts, allowing user behavior data to be transformed into training data suitable for LLM instruction tuning. The prompt consists of the following five parts:
@@
Task definition: A brief description of the next behavior prediction task, e.g., predict the next behavior the user will do from the candidate.
@@
Role-playing instruction: This defines the expected role and task for the LLM to assume and complete (Shanahan et al., 2023), e.g., you are a smart user’s mobile phone assistant, which can infer the user’s mobile phone behavior preferences based on the historical behavior history of the user.
@@
Historical behavior sequence (Input): Each element in is replaced with a textual description, e.g.,(1,16,home,Exercise), …,(1,20,home,Online shopping),(1,20,home,Video).
@@
Candidate set (Input): A collection of all possible behaviors the LLM can predict, constraining the range of behaviors to choose from, e.g., Weather Check, Music,…, Cycling, Reading.
@@
Instruction Fine-Tuning LLMs. To quickly adapt an LLM to behavior prediction tasks, we adopt instruction fine-tuning techniques (Ouyang et al., 2022). Based on the above prompt design, we transform the original user behavior dataset into a text-based instruction dataset , where and represent the corresponding natural language form of the input and output. The LLM is then optimized using the following next-token prediction loss:
where represents the learnable parameters of the LLM, refers to the -th token of , and indicates all tokens preceding . Moreover, to efficiently fine-tune the LLM with its vast number of parameters, we adopt a well-known parameter-efficient fine-tuning technique, i.e., LoRA (Hu et al., 2021). LoRA freezes the original parameter weights, decomposing them into the original part (which remains frozen) and an additional low-rank matrix , which can be updated much more efficiently while maintaining the performance of the LLM. By framing the task of predicting the next user behavior as predicting the next language token, we can more effectively leverage the LLM’s knowledge from its pretraining stage.
3.1.2. Progressive fine-tuning with behavior data
@@
As shown in Figure 1(a), user behavior data exhibits a significant class imbalance issue, where many types of behaviors occur infrequently in daily life, resulting in a long-tailed distribution. When transforming the task of predicting a user’s next behavior into generating the corresponding language token, the fine-tuned LLM still suffers from this class imbalance, leading to poor prediction performance for long-tailed behaviors, as demonstrated by “Llama3.1-8B-FT” in Figure 1(b). Motivated by the discovery that an LLM fine-tuned on a subset of frequently occurring behaviors (anchor behaviors) can surprisingly serve as an effective zero-shot predictor for other, less frequent behaviors (tail behaviors), performing even better than fine-tuning on the full set of behavioral data, we design a progressive fine-tuning strategy to enhance BehaviorLM’s predictive capability across diverse user behaviors.
@@
The proposed fine-tuning strategy consists of two progressive stages (Figure 2). In the first stage, the LLM is fine-tuned using anchor behavior data to specialize in user behavior prediction while retaining its inherent rich knowledge of long-tailed behaviors. In the second stage, the model is further fine-tuned on a class-balanced dataset covering all behaviors in a few-shot manner, enhancing its predictive capability for tail behaviors. The first stage helps the LLM become a specialist in predicting anchor behaviors, while the second stage transforms it into a generalist capable of predicting both anchor and tail behaviors.
3.2. Multitask-Enhanced Anchor Behavior Tuning (A-Tuning)